# OLMoE[Uncaptioned image]: Open Mixture-of-Experts Language Models

### v1

@@ Abstract
We introduce OLMoE, a fully open, state-of-the-art language model leveraging sparse Mixture-of-Experts (MoE). OLMoE-1B-7B has 7 billion (B) parameters but uses only 1B per input token. We pretrain it on 5 trillion tokens and further adapt it to create OLMoE-1B-7B-Instruct. Our models outperform all available models with similar active parameters, even surpassing larger ones like Llama2-13B-Chat and DeepSeekMoE-16B. We present various experiments on MoE training, analyze routing in our model showing high specialization, and open-source all aspects of our work: model weights, training data, code, and logs.
[Uncaptioned image] Weights https://hf.co/allenai/OLMoE-1B-7B-0924
Data {https://hf.co/datasets/allenai/OLMoE-mix-0924}
[Uncaptioned image] Code https://github.com/allenai/OLMoE
[Uncaptioned image] Logs https://wandb.ai/ai2-llm/olmoe/reports/
@@ 1 Introduction
Most MoE models, however, are closed-source: while some have publicly released model weights [43, 78, 156, 176, 178], they offer limited to no information about their training data, code, or recipes (see Figure 1). While there have been prior efforts in making language modeling research fully accessible [18, 64, 88, 102, 192, 208], they have been largely limited to dense LMs. This comes despite MoEs requiring more openness as they add complex new design questions to LMs, such as how many total versus active parameters to use, whether to use many small or few large experts, if experts should be shared, and what routing algorithm to use. The lack of open resources and findings about these details prevents the field from building cost-efficient open MoEs that approach the capabilities of closed-source frontier models.
@@ 2 Pretraining and Adaptation
OLMoE is a decoder-only LM consisting of $N_{L}$ transformer [183] layers. The feedforward network (FFN) in dense models like OLMo [64], is replaced with an MoE module consisting of $N_{E}$ smaller FFN modules called experts, of which a subset of $k$ experts are activated for each processed input token $x$ (also see Figure 2):
MoE $\text{MoE module}(x)=\sum_{i=1}^{\text{Top}-k(r(x))}\mathrm{softmax}\left(r(x)\right)_{i}{E_{i}(x)}$ (1)
@@
We use a mix of data from DCLM [89] and Dolma 1.7 [161], which includes the following: (1) a quality-filtered subset of Common Crawl, referred to as DCLM-Baseline, (2) StarCoder, Algebraic Stack and arXiv, used in both DCLM and Dolma 1.7, and (3) peS2o and Wikipedia from Dolma 1.7. We refer to our pretraining dataset as OLMoE-Mix.
To all sources above, we apply a filter that removes all documents with a sequence of 32 or more repeated n-grams, where an n-gram is any span of 1 to 13 tokens. For the StarCoder subset, we also remove any document that is either from a repository with fewer than 2 stars on GitHub, or whose most frequent word constitutes over 30% of the document, or whose top-2 most frequent words constitute over 50% of the document.
@@
We create OLMoE-1B-7B-Instruct by following a standard adaptation recipe split into instruction tuning [117, 189, 147, 154, 205] followed by preference tuning [31, 15, 136, 54] building on prior open models [182, 75, 186]. In our instruction tuning dataset, we add more code and math data to boost performance on downstream coding and math applications. Other models, such as GPT-4 [126] and Llama 3 [50], instead include samples from math datasets like GSM8k [35] or MATH [70] during pretraining. We also include No Robots and a subset of Daring Anteater as they are of high quality and add diversity, two key factors for successful adaptation [186, 215, 103, 119]. We describe our adaptation datasets in Table 3 and hyperparameters in Appendix B.
@@ 3 Results
OpenMoE-3B-9B [198] 2.9B [Uncaptioned image] 27.4 44.4 29.3 50.6 63.3 51.9
@@
Table 4: OLMoE-1B-7B after pretraining versus larger MoEs and dense LMs. We compare with dense LMs that assimilate OLMoE-1B-7B either in active parameters (1B, approximates speed and cost) or total parameters (7B, approximates memory requirements). Model names contain rounded parameter counts: model-active-total for MoEs and model-total for dense LMs (this leads to some differences to official names, e.g., while called “Gemma2-2B” it actually has 2.6B active and total parameters [175]). Chall. = Challenge. We run all evaluations ourselves with 5 few-shots, see Appendix C for details.
@@
Table 5: OLMoE-1B-7B after adaptation versus other models. We find the JetMoE chat model (https://hf.co/jetmoe/jetmoe-8b-chat) is broken leading to random scores thus we exclude it. Names contain rounded parameter counts: model-active-total for MoEs and model-total for dense LMs. We run all evaluations ourselves (Appendix C). Models use different mixes for adaptation, e.g., OLMoE is trained on an improved version of the pipeline used for OLMo models.
@@ 6 Related Work
A variety of model families have been proposed under varying degrees of openness commonly categorized based on whether model weights are available. Closed-weight models include GPT [24, 126], Gemini [172, 173], PaLM [30, 9], Reka [179], and open-weight ones include Llama [180, 181, 50], Mistral [77, 78], Gemma [174, 175], Falcon [8, 130], MPT [177], Qwen [13, 200], GLM [61], Yi [2], DeepSeek [42, 43, 39], Nemotron [128, 125, 188], InternLM [26], Baichuan [199], Phi [67, 93, 1], StableLM [16], OPT [211]. However, besides model weights, training data and code are key to enabling scientific research of these models [104, 105] and distributing their benefits broadly [23]. There have been few releases also including data and code in addition to model weights which we refer to as “fully open-source”: BLOOM [192, 149, 122, 202], GPT-NeoX [21, 22, 184], StarCoder [91, 108, 5, 119, 219], Pythia [18], OLMo [64], LLM360 [102], Cerebras-GPT [46], DCLM [89], MAP-Neo [208], RWKV [131, 132], and SmolLM [6]. For Mixture-of-Experts only OpenMoE [198] aims to be fully open-source, however, its poor performance limits its usefulness. We release OLMoE-1B-7B as the first state-of-the-art Mixture-of-Experts LM that is fully open-source: model weights, data, code, and logs.
@@ 7 Conclusion
We open-source OLMoE-1B-7B and OLMoE-1B-7B-Instruct including model, data, code, and logs. At 1B active and 7B total parameters, our models yield state-of-the-art performance among models with a similar amount of active parameters even outperforming larger models including DeepSeekMoE-16B and Llama2-13B-Chat. We share various training experiments and define and analyze router saturation, expert co-activation, domain and vocabulary specialization of our model. Through our fully open release, we seek to help the field build better MoEs. We are excited about new iterations of OLMoE to close the gap between frontier models and fully open models.
@@
Shane Arora uploaded the models and helped with code review.
@@
Alexander Wettig analyzed load balancing, routing, and domain specialization, and helped edit the paper.
Binyuan Hui advised on pretraining.
@@ Acknowledgements
OLMoE would not be possible without the support of many individuals and institutions. We thank our teammates at the Allen Institute for AI, Contextual AI, and the University of Washington for their support, especially Aditya Kusupati, Ananya Harsh Jha, Caitlin Wittlif, Carissa Schoenick, Costa Huang, Crystal Nam, David Atkinson, Emma Strubell, Faeze Brahman, Hamish Ivison, Karel D’Oosterlinck, Matt Latzke, Ian Magnusson, Jack Merullo, Jay Chen, Jennifer Dumas, Jiacheng Liu, Johann Dahm, Luke Zettlemoyer, Michael Schmitz, Michael Wilson, Pradeep Dasigi, Sahil Verma, Sam Skjonsberg, Sophie Lebrecht, Stas Bekman, Taira Anderson, Valentina Pyatkin, Yanai Elazar, Yizhong Wang, and Yoganand Chandrasekhar. We also thank Armen Aghajanyan, Akshat Shrivastava, Colin Raffel, Haokun Liu, Ludwig Schmidt, and Shayne Longpre. PWK is supported by the Singapore National Research Foundation and the National AI Group in the Singapore Ministry of Digital Development and Innovation under the AI Visiting Professorship Programme (award number AIVP-2024-001).
@@ Appendix B Training Configuration
We display the pretraining hyperparameter configuration of OLMoE-1B-7B in Appendix B comparing with other relevant models. We follow Groeneveld et al. [64] using the AdamW optimizer [106] with ZeRO [140] via PyTorch FSDP [212] and mixed-precision training [115]. Our main model settings differing from Groeneveld et al. [64] are: (1) MoE-related changes: OLMoE-1B-7B is a sparsely activated decoder-only transformer [183] using dropless Mixture-of-Experts [58]. Unlike most prior MoEs, we use a high granularity [39, 85] with 64 small experts with an FFN dimension of just 1,024 rather than a few large experts. We further use two auxiliary losses: router z-loss [220] and load balancing loss [152]. (2) Stability improvements: (a) We use a truncated normal initialization with a standard deviation of 0.02 and a minimum (maximum) cut-off of -0.06 (0.06) corresponding to three standard deviations. (b) We use QK normalization [171, 112, 44]. (c) We use RMSNorm [207] instead of the non-parametric LayerNorm used in Groeneveld et al. [64]. (3) Performance improvements: Besides some of the stability improvements which also impact performance, we also reduce the AdamW epsilon to 1.0E-08 from the 1.0E-05 used in Groeneveld et al. [64] to speed up convergence. Finally, we train OLMoE-1B-7B for significantly longer than all prior OLMo models amounting to 5T tokens and thus more than one epoch (1.2) following Muennighoff et al. [120]. We shuffle the pretraining dataset before starting the second epoch. For the final 100B tokens, we decay the learning rate linearly from 5.0E-04 to 0 (“annealing”). We experiment with many of these settings in §4.
@@ Appendix C Evaluation Setup
ARC-C [34] CF 0 token val max(MCF,CF) 5 pmi test
ARC-E [34] CF 0 none val max(MCF,CF) 5 character test
@@
CSQA [168] CF 0 token val max(MCF,CF) 5 pmi val
HellaSwag [206] CF 0 token val max(MCF,CF) 5 character val
MMLU [69] MCF 5 none val max(MCF,CF) 5 character test
MMLU Var CF 0-5 token val - - - -
OBQA [116] CF 0 token val max(MCF,CF) 5 pmi test
PIQA [20] CF 0 token val max(MCF,CF) 5 character val
@@
SocialIQA [148] CF 0 token val max(MCF,CF) 5 character val
@@
Table 11: Summary of downstream evaluation during and after pretraining (OLMES). ARC-C and ARC-E refer to ARC-Challenge and -Easy, CSQA=CommonsenseQA, OBQA=OpenBookQA, CF=Completion/Cloze formulation, MCF=Multiple-choice formulation, pmi=pointwise-mutual-information, Var=variants referring to the use of few-shots varying from 0-5.
@@
We evaluate using a similar in-loop evaluation setup as Groeneveld et al. [64], with the addition of more tasks such as CommonsenseQA, PIQA, and different implementations of MMLU. Following Groeneveld et al. [64], for the majority of the tasks, we perform 0-shot evaluation using the Completion/Cloze formulation (CF), ranking each answer string using language model probabilities. In terms of probability normalization, there is either no normalization (none) or normalization by the number of tokens in the answer (token) when ranking solely based on probability may heavily favor shorter answers [24]. For MMLU, the in-loop evaluation also includes a setup where we increase the total number of instances by including a range of 0-shot to 5-shot setups together as we found this provides smoother trends as the training proceeds (“MMLU Var”). We also include the Multiple-choice formulation (MCF) version of MMLU, scoring prediction of answer labels like A/B/C/D, which generally starts to rise only later in training as models only gain the multiple-choice capability later (at around 1T tokens for OLMoE-1B-7B in Figure 25). We also evaluate perplexity on selected validation sets from Paloma [110, 142, 59, 161, 95, 114]. All code used for evaluation during pretraining is at https://github.com/allenai/OLMo/tree/61ac104d616ec5435db225796e5c7532c9abd95a/olmo/eval.
@@ Appendix E Additional Evaluation
Mistral-7B $78.6^{\dagger}$ $90.8^{\dagger}$ 89.3 $72.4^{\dagger}$ 83.0 $64.0^{\dagger}$ $80.6^{\dagger}$ 82.8 $71.3^{\dagger}$ 77.9 79.1
OLMo-7B (0724) $68.0^{\dagger}$ $85.7^{\dagger}$ 85.3 $85.4^{\dagger}$ 80.5 $54.9^{\dagger}$ $67.6^{\dagger}$ 79.3 $76.1^{\dagger}$ 73.2 75.6
DCLM-7B $79.8^{\dagger}$ $92.3^{\dagger}$ 87.0 77.0 82.3 $64.4^{\dagger}$ $79.6^{\dagger}$ 80.1 $71.2^{\dagger}$ 77.3 79.1
Llama2-7B 54.2 84.0 86.1 74.2 78.9 $46.2^{\dagger}$ 57.8 77.5 59.6 71.7 69.0
Llama3.1-8B $79.5^{\dagger}$ $91.7^{\dagger}$ 88.5 $74.3^{\dagger}$ 81.6 $66.9^{\dagger}$ $78.6^{\dagger}$ 81.1 $71.4^{\dagger}$ 76.6 79.0
Gemma2-9B $89.5^{\dagger}$ $95.5^{\dagger}$ 89.4 $78.8^{\dagger}$ $87.3^{\dagger}$ $70.6^{\dagger}$ $88.4^{\dagger}$ $86.1^{\dagger}$ $76.0^{\dagger}$ 78.8 84.0
@@
StableLM-2B $50.6^{\dagger}$ 75.3 82.3 $70.4^{\dagger}$ 70.3 $40.4^{\dagger}$ $56.6^{\dagger}$ 75.6 $64.3^{\dagger}$ 65.8 65.1
Gemma2-3B $67.5^{\dagger}$ $84.3^{\dagger}$ 83.6 $66.4^{\dagger}$ 74.6 $53.3^{\dagger}$ $68.8^{\dagger}$ 78.5 $64.7^{\dagger}$ 71.8 71.4
JetMoE-2B-9B $61.4^{\dagger}$ $81.9^{\dagger}$ 85.7 $75.3^{\dagger}$ 81.7 $49.1^{\dagger}$ $68.0^{\dagger}$ 80.3 $71.3^{\dagger}$ 70.7 72.5
OpenMoE-3B-9B 29.3 50.6 63.2 21.5 44.4 27.4 34.6 63.3 42.9 $51.9^{\dagger}$ 42.9
DeepSeek-3B-16B 53.4 82.7 81.9 72.7 80.4 $45.5^{\dagger}$ 58.4 80.1 59.9 73.2 68.8
Qwen1.5-3B-14B $77.4^{\dagger}$ $91.6^{\dagger}$ 85.0 $81.4^{\dagger}$ 80.0 $62.4^{\dagger}$ $80.6^{\dagger}$ 81.0 $74.1^{\dagger}$ 72.3 78.6
@@
Pythia-1B 31.4 63.4 $56.8^{\dagger}$ 50.9 48.0 31.1 40.4 68.9 46.4 52.7 49.0
DCLM-1B $57.6^{\dagger}$ 79.5 80.9 71.3 75.1 $48.5^{\dagger}$ $60.0^{\dagger}$ 76.6 $60.5^{\dagger}$ 68.1 67.8
OLMoE-1B-7B $62.1^{\dagger}$ 84.2 79.2 72.9 80.0 $54.1^{\dagger}$ $65.4^{\dagger}$ 79.8 $63.0^{\dagger}$ 70.2 71.1
Table 12: More results on OLMES. indicates use of the MCF score, see Appendix C.

### v2

@@ Abstract
We introduce OLMoE, [^1] This paper describes the first OLMoE from 09/2024. See Appendix I for an overview of a newer version. a fully open, state-of-the-art language model leveraging sparse Mixture-of-Experts (MoE). OLMoE-1B-7B has 7 billion (B) parameters but uses only 1B per input token. We pretrain it on 5 trillion tokens and further adapt it to create OLMoE-1B-7B-Instruct. Our models outperform all available models with similar active parameters, even surpassing larger ones like Llama2-13B-Chat and DeepSeekMoE-16B. We present various experiments on MoE training, analyze routing in our model showing high specialization, and open-source all aspects of our work: model weights, training data, code, and logs.
[Uncaptioned image] Model hf.co/allenai/OLMoE-1B-7B-0924
Data hf.co/datasets/allenai/OLMoE-mix-0924
[Uncaptioned image] Code github.com/allenai/OLMoE
[Uncaptioned image] Logs wandb.ai/ai2-llm/olmoe/reports/
@@ 1 Introduction
Most MoE models, however, are closed-source: While some have publicly released model weights [43, 79, 158, 178, 180], they offer limited to no information about their training data, code, or recipes (see Figure 1). While there have been prior efforts to make language modeling research fully accessible [18, 65, 90, 103, 193, 209], they have been largely limited to dense LMs. This comes despite MoEs requiring more openness as they add complex new design questions to LMs, such as how many total versus active parameters to use, whether to use many small or few large experts, if experts should be shared, and what routing algorithm to use. The lack of open resources and findings about these details prevents the field from building cost-efficient open MoEs that approach the capabilities of closed-source frontier models.
@@ 2 Pretraining and Adaptation
OLMoE is a decoder-only LM consisting of $N_{L}$ transformer [185] layers. The feedforward network (FFN) in dense models like OLMo [65], is replaced with an MoE module consisting of $N_{E}$ smaller FFN modules called experts, of which a subset of $k$ experts is activated for each processed input token $x$ (also see Figure 2):
MoE $\text{MoE module}(x)=\sum_{i\in\text{Top}-k(r(x))}\mathrm{softmax}\left(r(x)\right)_{i}{E_{i}(x)}$ (1)
@@
We mix data from DCLM [90] and Dolma 1.7 [163], which includes: (1) a quality-filtered subset of Common Crawl, referred to as DCLM-Baseline, (2) StarCoder, Algebraic Stack and arXiv, used in both DCLM and Dolma 1.7, and (3) peS2o and Wikipedia from Dolma 1.7. We refer to our pretraining dataset as OLMoE-Mix.
To all sources above, we apply a filter that removes all documents with a sequence of 32 or more repeated n-grams, where an n-gram is any span of 1 to 13 tokens. For the StarCoder subset, we also remove any document from a repository with fewer than 2 stars on GitHub, whose most frequent word constitutes over 30% of the document, or whose top-2 most frequent words constitute over 50% of the document.
@@
We create OLMoE-1B-7B-Instruct by following a standard adaptation recipe split into instruction tuning [118, 190, 149, 156, 206] followed by preference tuning [31, 15, 138, 54] building on prior open models [184, 76, 188]. In our instruction tuning dataset, we add more code and math data to boost performance on downstream coding and math applications. Other models, such as GPT-4 [128] and Llama 3 [50] similarly include samples from math datasets like GSM8k [35] or MATH [71] during pretraining. We also include No Robots and a subset of Daring Anteater as they are of high quality and add diversity, two key factors for successful adaptation [188, 216, 104, 120]. We describe our adaptation datasets in Table 3 and hyperparameters in Appendix B.
@@ 3 Results
OpenMoE-3B-9B [199] 2.6B [Uncaptioned image] 27.4 44.4 29.3 50.6 63.3 51.9
@@
Table 4: OLMoE-1B-7B after pretraining versus larger MoEs and dense LMs. We compare with dense LMs close to OLMoE-1B-7B either in active parameters (1B, approximates speed and cost) or total parameters (7B, approximates memory requirements). Model names contain rounded parameter counts: model-active-total for MoEs and model-total for dense LMs (this leads to some differences to official names, e.g., while called “Gemma2-2B” it actually has 2.6B active and total parameters [177]). Chall. = Challenge. We run all evaluations ourselves with 5 few-shots, see Appendix C for details.
@@
Table 5: OLMoE-1B-7B after adaptation versus other models. We find the JetMoE chat model (https://hf.co/jetmoe/jetmoe-8b-chat) has random scores thus we exclude it. Model names contain rounded parameter counts: model-active-total for MoEs and model-total for dense LMs. We run all evaluations ourselves (Appendix C). Models use different mixes for adaptation, e.g., OLMoE is trained on an improved version of the pipeline used for OLMo models.
@@ 6 Related Work
A variety of model families have been proposed under varying degrees of openness commonly categorized based on whether model weights are available. Closed-weight models include GPT [24, 128], Gemini [174, 175], PaLM [30, 9], Reka [181], and open-weight ones include Llama [182, 183, 50], Mistral [78, 79], Gemma [176, 177], Falcon [8, 132], MPT [179], Qwen [13, 201], GLM [61], Yi [2], DeepSeek [42, 43, 39], Nemotron [130, 126], Zamba [62], InternLM [26], Baichuan [200], Phi [68, 94, 1], StableLM [16], OPT [212]. However, besides model weights, training data and code are key to enabling scientific research of these models [105, 106] and distributing their benefits broadly [23]. There have been few releases also including data and code in addition to model weights which we refer to as “fully open-source”: BLOOM [193, 151, 123, 203], GPT-NeoX [21, 22, 186], StarCoder [92, 109, 5, 120, 220], Pythia [18], OLMo [65], LLM360 [103], Cerebras-GPT [46], DCLM [90], MAP-Neo [209], RWKV [133, 134], and SmolLM [6]. For Mixture-of-Experts only OpenMoE [199] aims to be fully open-source, however, its poor performance limits its usefulness. We release OLMoE-1B-7B as the first state-of-the-art Mixture-of-Experts LM that is fully open-source: model weights, data, code, and logs.
@@ 7 Conclusion
We open-source OLMoE-1B-7B and OLMoE-1B-7B-Instruct including model, data, code, and logs. At 1B active and 7B total parameters, our models yield state-of-the-art performance among models with a similar amount of active parameters even outperforming larger models including DeepSeekMoE-16B and Llama2-13B-Chat. We share various training experiments and define and analyze router saturation, expert co-activation, domain and vocabulary specialization of our model. Through our fully open release, we seek to help the field build better MoEs. We are excited about more iterations of OLMoE to close the gap between frontier models and fully open models.
@@
Shane Arora uploaded the models, helped with code review and framework integration.
@@
Alexander Wettig advised on pretraining, analyzed load balancing, routing, and domain specialization, and helped edit the paper.
Binyuan Hui advised on pretraining and helped with plotting and framework integration.
@@ Acknowledgements
OLMoE would not be possible without the support of many individuals and institutions. We thank our teammates at the Allen Institute for AI, Contextual AI, and the University of Washington for their support, especially Aditya Kusupati, Ananya Harsh Jha, Caitlin Wittlif, Carissa Schoenick, Costa Huang, Crystal Nam, David Atkinson, Emma Strubell, Faeze Brahman, Hamish Ivison, Karel D’Oosterlinck, Matt Latzke, Ian Magnusson, Jack Merullo, Jay Chen, Jennifer Dumas, Jiacheng Liu, Johann Dahm, Luke Zettlemoyer, Michael Schmitz, Michael Wilson, Pradeep Dasigi, Sahil Verma, Sam Skjonsberg, Sophie Lebrecht, Stas Bekman, Taira Anderson, Valentina Pyatkin, Yanai Elazar, Yizhong Wang, and Yoganand Chandrasekhar. We also thank Armen Aghajanyan, Akshat Shrivastava, Colin Raffel, Haokun Liu, Ludwig Schmidt, Mengzhou Xia, Shayne Longpre, Sheng Shen, and Zexuan Zhong. PWK is supported by the Singapore National Research Foundation and the National AI Group in the Singapore Ministry of Digital Development and Innovation under the AI Visiting Professorship Programme (award number AIVP-2024-001).
@@ Appendix B Training Configuration
We display the pretraining hyperparameter configuration of OLMoE-1B-7B in Appendix B comparing with other relevant models. We follow Groeneveld et al. [65] using the AdamW optimizer [107] with ZeRO [142] via PyTorch FSDP [213] and mixed-precision training [116]. Our main model settings differing from Groeneveld et al. [65] are: (1) MoE-related changes: OLMoE-1B-7B is a sparsely activated decoder-only transformer [185] using dropless Mixture-of-Experts [58]. Unlike most prior MoEs, we use a high granularity [39, 86] with 64 small experts with an FFN dimension of just 1,024 rather than a few large experts. We further use two auxiliary losses: router z-loss [221] and load balancing loss [154]. (2) Stability improvements: (a) We use a truncated normal initialization with a standard deviation of 0.02 and a minimum (maximum) cut-off of -0.06 (0.06) corresponding to three standard deviations. (b) We use QK normalization [173, 113, 44]. (c) We use RMSNorm [208] instead of the non-parametric LayerNorm used in Groeneveld et al. [65]. (3) Performance improvements: Besides some of the stability improvements which also impact performance, we also reduce the AdamW epsilon to 1.0E-08 from the 1.0E-05 used in Groeneveld et al. [65] to speed up convergence. Finally, we train OLMoE-1B-7B for significantly longer than all prior OLMo models amounting to 5T tokens and thus more than one epoch (1.3) following Muennighoff et al. [121]. We shuffle the pretraining dataset before starting the second epoch. For the final 100B tokens, we decay the learning rate linearly from 5.0E-04 to 0. We experiment with many of these settings in §4.
@@ Appendix C Evaluation Setup
ARC-C [34] CF 0 char val max(MCF,CF) 5 pmi test
ARC-E [34] CF 0 none val max(MCF,CF) 5 char test
@@
CSQA [170] CF 0 char val max(MCF,CF) 5 pmi val
HellaSwag [207] CF 0 char val max(MCF,CF) 5 char val
MMLU [70] MCF 5 none val max(MCF,CF) 5 char test
MMLU Var CF 0-5 char val - - - -
OBQA [117] CF 0 char val max(MCF,CF) 5 pmi test
PIQA [20] CF 0 char val max(MCF,CF) 5 char val
@@
SocialIQA [150] CF 0 char val max(MCF,CF) 5 char val
@@
Table 11: Summary of downstream evaluation during and after pretraining (OLMES). ARC-C and ARC-E refer to ARC-Challenge and -Easy, CSQA=CommonsenseQA, OBQA=OpenBookQA, CF=Completion/Cloze formulation, MCF=Multiple-choice formulation, pmi=pointwise-mutual-information, char=per-character, Var=variants referring to the use of few-shots varying from 0-5.
@@
We evaluate using a similar in-loop evaluation setup as Groeneveld et al. [65], with the addition of more tasks such as CommonsenseQA, PIQA, and different implementations of MMLU. Following Groeneveld et al. [65], for the majority of the tasks, we perform 0-shot evaluation using the Completion/Cloze Formulation (CF), ranking each answer string using language model probabilities. In terms of probability normalization, there is either no normalization (none) or normalization by the number of characters in the answer (char) when ranking solely based on probability may heavily favor shorter answers [24]. For MMLU, the in-loop evaluation also includes a setup where we increase the total number of instances by including a range of 0-shot to 5-shot setups together as we found this provides smoother trends as the training proceeds (“MMLU Var”). We also include the Multiple-Choice Formulation (MCF) version of MMLU, scoring prediction of answer labels like A/B/C/D, which generally starts to rise only later in training as models only gain the multiple-choice capability later (at around 1T tokens for OLMoE-1B-7B in Figure 25). We also evaluate perplexity on selected validation sets from Paloma [111, 144, 59, 163, 96, 115]. All code used for evaluation during pretraining is at https://github.com/allenai/OLMo/tree/61ac104d616ec5435db225796e5c7532c9abd95a/olmo/eval.
@@ Appendix E Additional Evaluation
Mistral-7B 78.6 90.8 89.3 72.4 83.0 64.0 80.6 82.8 71.3 77.9 79.1
OLMo-7B (0724) 68.0 85.7 85.3 85.4 80.5 54.9 67.6 79.3 76.1 73.2 75.6
DCLM-7B 79.8 92.3 87.0 77.0 82.3 64.4 79.6 80.1 $71.2^{\dagger}$ 77.3 79.1
Llama2-7B 54.2 84.0 86.1 74.2 78.9 46.2 57.8 77.5 59.6 71.7 69.0
Llama3.1-8B 79.5 91.7 88.5 74.3 81.6 66.9 78.6 81.1 71.4 76.6 79.0
Gemma2-9B 89.5 95.5 89.4 78.8 87.3 70.6 88.4 86.1 76.0 78.8 84.0
@@
StableLM-2B 50.6 75.3 82.3 70.4 70.3 40.4 56.6 75.6 64.3 65.8 65.1
Gemma2-3B 67.5 84.3 83.6 66.4 74.6 53.3 68.8 78.5 64.7 71.8 71.4
JetMoE-2B-9B 61.4 81.9 85.7 75.3 81.7 49.1 68.0 80.3 71.3 70.7 72.5
OpenMoE-3B-9B 29.3 50.6 63.2 21.5 44.4 27.4 34.6 63.3 42.9 51.9 42.9
DeepSeek-3B-16B 53.4 82.7 81.9 72.7 80.4 45.5 58.4 80.1 59.9 73.2 68.8
DeepSeekV2-2B-16B 74.0 88.9 84.7 73.8 81.9 58.8 72.4 80.2 69.1 74.0 75.8
Llama3.2-3B 69.6 85.1 78.3 69.0 77.0 57.8 67.2 77.4 64.9 69.9 71.6
Qwen1.5-3B-14B 77.4 91.6 85.0 81.4 80.0 $62.4^{\dagger}$ 80.6 81.0 74.1 72.3 78.6
@@
Pythia-1B 31.4 63.4 56.8 50.9 48.0 31.1 40.4 68.9 46.4 52.7 49.0
Llama3.2-1B 43.5 71.6 69.4 59.6 67.3 38.2 42.0 73.7 52.0 62.5 58.0
Zamba2-1B 55.0 85.4 76.1 70.1 73.4 44.73 59.8 76.6 58.4 67.2 66.7
DCLM-1B 57.6 79.5 80.9 71.3 75.1 48.5 60.0 76.6 60.5 68.1 67.8
OLMoE-1B-7B 62.1 84.2 79.2 72.9 80.0 54.1 65.4 79.8 63.0 70.2 71.1
Table 12: More results on OLMES. indicates use of the MCF score, see Appendix C. See Table 4 for details on naming and a summary of these results.